Thirty-six percent of primary care doctors think their patients receive unnecessary medical care. If required, round your answer to four decimal places.

(a) Suppose a sample of 300 primary care doctors was taken. Show the distribution of the sample proportion of doctors who think their patients receive unnecessary medical care.
np =
n(1-p) =
E(p) =
σ(p) =
(b) Suppose a sample of 500 primary care doctors was taken. Show the distribution of the sample proportion of doctors who think their patients receive unnecessary medical care.
np =
n(1-p) =
E(p) =
σ(p) =
(c) Suppose a sample of 1,000 primary care doctors was taken. Show the distribution of the sample proportion of doctors who think their patients receive unnecessary medical care.
np =
n(1-p) =
E(p) =
σ(p) =
(d) In which of the preceding three cases, part (a) or part (b) or part (c), is the standard error of p smallest? Why?
The standard error is the smallest in - Select your answer -part (a)part (b)part (c)Item 13 because p is - Select your answer -the samenot the sameItem 14 in parts (a), (b), and (c) and the sample size is largest in - Select your answer -part (a)part (b)part (c)Item 15 .
Answer #1

c) Given n=1000 and p=0.36 np=1000*0.36=360 n(1-p=1000*(1-0.36)=640 E(p)=p=0.36 p(1-P) _ O(P) = 0.36(1-0.36) =0,015179 1000 d